If a and b are odd integers,which of the following must also be an odd integer?
(i) (a+1)b,(ii)(a+1)+b(iii)(a+1)-b.

Answer by stanbon(75887)

a and b are odd integers,which of the following must also be an odd integer?
(i) (a+1)b = even*odd = even
-------
(ii)(a+1)+b = even + odd = odd
------------
(iii)(a+1)-b = even - odd = odd
----------------
Cheers,
Stan H.
--------------